40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 235 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  5705   Wed Oct 19 18:16:53 2011 JenneUpdatePSLPMC found unlocked

I just relocked the PMC.  I don't know why it was unlocked.

  5704   Wed Oct 19 17:33:07 2011 kiwamuUpdateGreen LockingETMY mechanical shutter : fully functional

The mechanical shutter on the Y end is now fully functional.

 

 It is newly named to 'C1AUX_GREEN_Y_Shutter' in the EPICS world.

It uses the same binary output channel which had been served for the POY shutter.

To change the EPICS name I edited a db file called ShutterInterLock.db, which resides in /cvs/cds/caltech/target/c1aux.

After editing the file I rebooted the c1aux machine, by telnet and the reboot command in order to make the change effective.

Also I added this new shutter on the ALS overview screen (see the attachment below)

MechShutter.png

Quote from #5701

The ETMY  shutter can be remotely switched from medm screen POY of mechanical shutters.

The new cable from ETMY controller goes to east vertex EV-ISCT where it is connected to POY shutter hook up BNC cable.

  5703   Wed Oct 19 17:21:16 2011 KojiSummaryLSCModification on the RFPD interface cards

I have modified all of the three RFPD interface cards to be enabled permanently.
This prevents an accidental disabling caused by a stray voltage of the logic input (or whatever),
which was reported in multiple occasions by Anamaria and me.

The logic ICs (74LS04) for buffering of the EPICS switches were removed by 14pin sockets with additional wires soldered.
The modification shorts the inputs to the second logic chips, resulting in the permanent enabling of the PD circuit.

  5702   Wed Oct 19 16:53:38 2011 kiwamuUpdateCDSsome screens need labels

Untitled.png

Some of the sub-suspension screens need labels to describe what those row and column are.

  5701   Wed Oct 19 16:28:35 2011 steveUpdateGreen LockingETMY mechanical shutter is working

Quote:

Uniblitz mechanical shutter installed in the green beam path at ETMY-ISCT  The remote control cable has not been connected.

The ETMY  shutter can be remotely switched from medm screen POY of mechanical shutters.

The new cable from ETMY controller goes to east vertex EV-ISCT where it is connected to POY shutter hook up BNC cable.

  5700   Wed Oct 19 15:48:20 2011 MirkoUpdatePEMMoved the STS1 from x-arm end to vortex

[Jenne, Mirko]

We moved our one STS1 from the x-arm end to the vortex. We record the data as STS1 in c1pem @ 256Hz. x is still north-south.

JD:  This is actually an STS-2.  We just call it C1:PEM-SEIS_STS1.... to differentiate the 3 STSs that we have from one another (assuming we plug in the other two).

19102011061.jpg

  5699   Wed Oct 19 15:46:49 2011 kiwamuUpdateSUStaking care of SRM

Quote from #5691

I am going to check the sensing matrix with the new free swinging spectra (#5690)

 

The SRM input matrix has been readjusted.

However still there is the unwanted coupling from the POS drive to SIDE signal and from the SIDE drive to POS signal.

      BADNESS
  SRM  SRM.png       pit     yaw     pos     side    butt
UL    0.871   0.975   1.115  -0.295   1.096  
UR    1.015  -1.025   1.128  -0.140  -1.053  
LR   -0.985  -0.984   0.885  -0.088   0.831  
LL   -1.129   1.016   0.872  -0.243  -1.020  
SD    0.084   0.061   3.534   1.000  -0.018  
 
 4.24965

 

  5698   Wed Oct 19 14:03:41 2011 kiwamuUpdateGeneralvent prep : dichroic mirrors

Status update on dichroic mirror:

 I got the specification sheet of an aLIGO 2" dichroic mirror from Lisa.

https://dcc.ligo.org/cgi-bin/private/DocDB/ShowDocument?docid=25232

This is the one from ATF. It has low loss, a high R for 1064 nm and high T for 532 nm. So it matches our needs.

Based on this sheet we may reset some of the parameters in the specification (e.g. incident angle and etc.,) and will get a quote from ATF.

We will buy 3 of them, including 1 spare. First I need to review the specification.

  5697   Wed Oct 19 13:45:11 2011 SureshUpdate40m UpgradingActive Tiptilts from LLO moved to clean shelf along X arm

I have moved the active tip tilts that we brought over from LLO to the Clean Bureau along the X arm (closest to the ETMX). There are two tip tilts and a pack of spare parts.  I am sure that the tip tilts are clean, packed in the clean room at LLO.  I am not sure whether the spares are clean.  I have kept them together for now.

We need to open one of the Tip tilt packages to be sure what we have got.

  5696   Wed Oct 19 12:25:58 2011 SureshUpdate40m UpgradingActive Tiptilts from LLO moved to clean shelf along X arm

I have moved the active tip tilts that we brought over from LLO to the Clean Bureau along the X arm (closest to the ETMX). There are two tip tilts and a pack of spare parts.

  5695   Wed Oct 19 12:06:26 2011 MirkoUpdateCDSIncluded the MC servo channel in CDS

[Jamie, Mirko]

Included the 'Servo' output from the D040180 in c1ioo, which I hoped would be a better measure of the MC length fluctuations. It goes into ADC6, labeled CH7 on the physical board.
Servo-output => C1:IOO-MC_SERVO. (Already present is Out1-output => C1:IOO-MC_F).
At low freq. the servo signal is about 4.5dB bigger. Both are recorded at 256Hz now which is the reason for the downward slope at about 100Hz.

MC_F_versus_MC_SERVO.png

Coh_MC_F_MC_SERVO.png

  5694   Wed Oct 19 10:49:35 2011 kiwamuUpdateSUSSRM oscillation removed

Quote:

The SRM oplevs were found to be oscillating because of a small phase margin.

I reduced the gains to the half of them. The peak which Steve found today maybe due to this oscillation.

Quote from #5630

The SRM bounce peak 18.33 Hz. Suresh helped me to retune filter through Foton, but we failed to remove it. 

 Kiwamu removed the 18.3 Hz ocsillation by turning down the oplev servo gain.

  5693   Wed Oct 19 09:44:10 2011 steveUpdateSUSETMX oplev power spectrum

Quote:

 ETMX oplev had 6 mm diameter beam on the qpd.  I relayed the beam path with 2 lenses  to get back  3 mm beam on the qpd

BRC 037  -100 Bi _concave lens and PCX 25  200 VIS do the job. Unfortunately the concave lens has the AR 1064.

 

 

 Coherence at 1 and 2-3 Hz only. He/Ne laser intensity noise is not an issue.

  5692   Wed Oct 19 08:34:16 2011 steveUpdateSUSSRM sus damping restored

Sorry Kiwamu, I realized too late that you were freeswigging. Hopefully 4 hrs was enough.

  5691   Wed Oct 19 05:15:44 2011 kiwamuUpdateSUStaking care of SRM

I made some efforts to fix the situation of SRM but it is still bad.

The POS motion wasn't well damped. Something is wrong either (maybe both) sensing part or actuation part.

I am going to check the sensing matrix with the new free swinging spectra (#5690)

 

(Symptom)

 When I was trying to lock SRMI I found that the fringes observed at the AS camera didn't show spatial higher order modes, which is good.

So I thought the SRM suspension became quiet, but it actually wasn't. Because the RMS monitor of the SRM OSEMs already went to about 30 counts.

At the same time the opelev error signals were well suppressed. It means some DOFs which were insensitive to the oplev were ringing up, namely POS and SIDE.

According to the LSC error signal and the ASDC signal, I believe that the POS was going wild (although I didn't check the OSEM spectra).
 

(Some efforts)

 + Readjusted the f2a filters (see the attachment).

 + Tried to eliminate a coupling between the POS and SIDE drives by tweaking the output matrix.

    => In order to eliminate the coupling from the POS drive to SIDE sensor, I had to put a comparable factor into an element.

     So it might be possible that the POS sensor was showing the SIDE signal and vice versa.

      In order to check it I left SRM free swinging (#5690).

Quote from #5684

The main reason why I couldn't lock DRMI was that the suspensions were touchy and especially the SRM suspension wasn't good.

  5690   Wed Oct 19 04:23:58 2011 kiwamuUpdateSUSSRM free swinging test

The following optics were kicked:
SRM
Wed Oct 19 04:22:53 PDT 2011
1003058589
 

  5689   Tue Oct 18 22:47:09 2011 SureshConfigurationIOOMC autolocker script edited to shutdown and restart WFS loops

Quote:

I found that the MC WFS had large offset control signals going to the MC SUS. Even though the input switch was off, the integrators were holding the offset.

I have disabled the ASCPIT outputs in the MC SUS. Suresh is going to fix the MC autolocker script to gracefully handle the OFF and ON and then test the script before resuming the WFS testing.

MCL data for OAF may be suspect from this morning.

 I have edited (uncommented existing commands)  the following scripts to enable WFS locking to come on when the MC is locked.

1) $scripts$/MC/autolockMCmain40m*

2) $scripts$/MC/mcup

3) $scripts$/MC/mcdown

4) $scripts$/MC/WFS/mcwfson

5) $scripts$/MC/WFS/mcwfsoff.

I have checked that the autolocker script switches off the mcwfs when mc loses lock and then switches it on after re-obtaining lock.

 

  5688   Tue Oct 18 21:19:18 2011 ranaConfigurationIOOWFS disabled in SUS

I found that the MC WFS had large offset control signals going to the MC SUS. Even though the input switch was off, the integrators were holding the offset.

I have disabled the ASCPIT outputs in the MC SUS. Suresh is going to fix the MC autolocker script to gracefully handle the OFF and ON and then test the script before resuming the WFS testing.

MCL data for OAF may be suspect from this morning.

  5687   Tue Oct 18 20:50:19 2011 SureshUpdateIOOC1IOO and WFS associated screens

In keeping with the current protocol,  I have started to move all the user-built medm screens associated with C1IOO into the $screens$/c1ioo/master/ directory. 

I then edited the menu button in the sitemap.adl to point to the screens in the ..c1ioo/master/ directory.  All the screens in $screens$/c1ioo/ directory have been backed up into bak/.  I plan to edit the c1ioo model soon and at that time I will delete all the screens in the $screens$/c1ioo directory and let only the automatically regenerated screens  stay there.   If there are broken links to user-built screens associated with c1ioo, please copy the relevant screen to the master/ directory and edit the path in the menus.

 

  5686   Tue Oct 18 15:20:03 2011 kiwamuSummaryIOORFAM plan

[Suresh / Koji / Rana / Kiwamu]

Last night we had a discussion about what we do for the RFAM issue. Here is the plan.

 

(PLAN)

  1. Build and install an RFAM monitor (a.k.a StochMon ) with a combination of a power splitter, band-pass-filters and Wenzel RMS detectors.

       => Some ordering has started (#5682). The Wenzel RMS detectors are already in hands.

  2. Install a temperature sensor on the EOM. And if possible install it with a new EOM resonant box.

      => make a wheatstone bridge circuit, whose voltage is modulated with a local oscillator at 100 Hz or so.

  3. Install a broadband RFPD to monitor the RFAMs and connect it to the StochMon network.

      => Koji's broadband PD or a commercial RFPD (e.g. Newfocus 1811 or similar)

  4. Measure the response of the amount of the RFAM versus the temperature of the EO crystal.

      => to see whether if stabilizing the temperature stabilizes the RFAM or not.

  5.  Measure the long-term behavior of the RFAM.

      => to estimate the worst amount of the RFAM and the time scale of its variation

  6. Decide which physical quantity we will stabilize, the temperature or the amount of the RFAM.

  7. Implement a digital servo to stabilize the RFAMs by feeding signals back to a heater

     => we need to install a heater on the EOM.

  8. In parallel to those actions, figure out how much offsets each LSC error signal will have due to the current amount of the RFAMs.

    => Optickle simulations.

  9. Set some criteria on the allowed amount of the RFAMs

    => With some given offsets in the LSC error signal, we investigate what kind of (bad) effects we will have.

  5685   Tue Oct 18 10:04:41 2011 KojiUpdateLSCREFL165 removed from the table

The original REFL165 had ~50MHz/A dependence on the DC photocurrent.
The resistr R21, which was 2670 Ohm contrary to the original drawing, was replaced to 532 Ohm
to increase the feedforward gain by factor of 5.

The resulting dependence is reduced to ~0.5MHz/A although it has Q reduction of ~20% at 6mA.

Some concerns:

These transfer functions were measured between TEST IN and RF OUT while the diode was illuminated with the white light from a light bulb.

There looks some thermal effect on the resonant freq. If the white light illumination is suddenly removed, the bias compensation
is immediately removed but the resonance takes some time (~min) to come back to the original freq.

I am afraid that the light bulb gave too much heat on the surrounding PCB and lead unnecesarily high level dependence of the resonant freq on the DC current.

Or, if this thermal effect comes from the power consumption on the diode itself, we need to characterize it for aLIGO.

In order to check this, we need a test with the 1064nm illumination on the diode in stead of the light bulb.

  5684   Tue Oct 18 04:04:27 2011 kiwamuUpdateLSCmeasurement of sensing matrix : touchy SRM

I made some attempts to measure the sensing matrix of the central part.

I could measure the matrix in the PRMI configuration but wasn't able to measure the matrix in the DRMI configuration.

   => I will report the result of the PRMI sensing matrix tomorrow.

The main reason why I couldn't lock DRMI was that the suspensions were touchy and especially the SRM suspension wasn't good.

Some impacts due to the feedback during the lock acquisition completely kicks SRM away. 

The watchdogs' RMS monitor on SRM easily rang up to more than 10 counts once the acquisition started.This is quite bad.

Also the stability of the PRMI lock was strongly depending on the gains of the PRM oplevs.

I guess I have to revisit the vertex suspensions more carefully (i.e. f2a coupling, actuator output matrix, damping gains, input matrices, oplev filters)

otherwise any LSC works in the vertex will be totally in vain.

  5683   Mon Oct 17 23:56:34 2011 SureshUpdateIOOMC WFS Integrators switched on and WFS_MASTER screen updated

[Rana, Suresh]

     To see if the loops will stay locked when the Integrators in the servo are switched on, we stayed with the same simple output matrix (just 1 or -1 elements) and switched on the FM1 on all WFS servo filter banks.  We monitored the time domain error signals to see if engaging the locks made the error signals go to zero.  Most of the error signals did go to zero even when an intentional offset was introduced into the MC pitch of the suspension.

      We need to include TestPoints just before the Input Servo Matrix so that we can monitor the error signals without being affected by the gain changes in the WFS_GAIN slider.   These are currently not present in the C1IOO model and the position of the WFS_GAIN also has to be shifted to the other side of the Input matrix.

      The C1IOO_WFS_MASTER screen has been changed to the new one.  This incorporates filter banks for the MC_TRANS_P and _Y channels.  The screen is not yet fully functional but I am working on it and I it will continue to improve it.

WFS_MASTER_screenshot_20111017.png

  5682   Mon Oct 17 23:28:32 2011 ranaUpdateElectronicsStochMon

To get to the bottom of the RFAM mystery, we've got to resurrect the StochMon to trend the RFAM after the IMC.

We will put an 1811 on the MC_TRANS or IP_POS beam (the 1811 has an input noise of 2.5 pW/rHz).

Then the Stochmon has an input pre-amp, some crappy filters, and then Wenzel RMS->DC converters. We will replace the hand-made filters with the following ones from Mini-Circuits which happen to match our modulation frequencies perfectly:

11 MHz     SBP-10.7+

55 MHz     SBP-60+

29.5 MHz   SBP-30+

  5681   Mon Oct 17 22:20:42 2011 KojiUpdateLSCREFL165 removed from the table

Quote:

REFL165 removed from the table for the C(V) test

The PD was returned on the table.

The C(V) compensation path was modified and the change of the resonant freq was cancelled.
A more precise analysis comes later.

  5680   Mon Oct 17 17:07:30 2011 steveUpdateSUSETMX oplev returning beam od 3 mm

 ETMX oplev had 6 mm diameter beam on the qpd.  I relayed the beam path with 2 lenses  to get back  3 mm beam on the qpd

BRC 037  -100 Bi _concave lens and PCX 25  200 VIS do the job. Unfortunately the concave lens has the AR 1064.

 

 

  5679   Mon Oct 17 14:26:22 2011 MirkoUpdateCDSSeismic BLRMS channels, new RMS calculation

Quote:

[Rana, Koji, Mirko]

We looked into the CDS RMS block c-code as described in Rolfs RCG app guide. Seems the block uses a first order LP filter with a corner freq. / time of 20k execution cycles. There are also some weird thersholds at +-2000counts in there.

I was looking into implementing a hand-made RMS block, by squaring, filtering, rooting. The new RMS (left) seems nicer than the old one (bottom right). Signal was 141counts sinus at 4Hz.

Filters used: Before squaring: 4th order butterworth BP at given freq. & (new) 6th order inverse Chebyshew 20dB at 0.9*lower BP freq. and 1.1*upper BP freq. => about 1dB at BP freq.

                       After squaring: 4th order butterworth LP @ 1Hz.

C1PEM execution time increased from about 20us to about 45us.

Made a new medm screen with the respective filters in place of the empty C1PEM_OVERVIEW. Should go onto the sitemap.

New_RMS_vs_old_RMS.png

Original RMS LP is slower than 0.1Hz, see below for single LP at 0.1Hz in the new RMS. Original RMS is faster than single LP @ 0.01Hz

Original_RMS_LP_slower_than_0.1Hz.png

Some of the channels are recorded as 256Hz DAQ channels now. Need to figure out how to record these as 16Hz EPICS channls.

 Channels are now going into EPICS channels (e.g. C1:PEM-ACC1_RMS_1_3 ). Adapted the PEM_SLOW.ini file. Channels don't yet show up in dataviewer. Probably due to other C1PEM maschine

  5678   Mon Oct 17 11:40:44 2011 KojiUpdateLSCREFL165 removed from the table

REFL165 removed from the table for the C(V) test

  5677   Mon Oct 17 11:06:31 2011 MirkoUpdateCDSPiping data from c1lsc to c1oaf

Changed, recompiled, installed and restarted c1rfm and c1oaf to get the MC1-3 Pitch and Yaw data into the c1oaf model.
Also changed c1oaf to use MCL as a witness channel (as well as an actuator).
Added the changes to svn.

  5676   Mon Oct 17 10:43:14 2011 MirkoUpdateCDSCommited changes to c1rfm

I want to make changes to c1rfm. Found uncommited changes to it from Sept 27. Since we recompiled it since then it should be safe to commit them, so I did. See svn log for details.

  5675   Mon Oct 17 07:57:24 2011 steveUpdateSUSETMX sus damping restored
  5674   Sun Oct 16 05:35:18 2011 ranaUpdateComputer Scripts / ProgramsFailing to set SUS summary screen values

Quote:

Quote:

I am trying to run Rana's setSensors.py script, but am failing.  Any inspiration would be appreciated:

rosalba:SUS_SUMMARY>./setSensors.py 1001708529 500 .1 .25
['./setSensors.py', '1001708529', '500', '.1', '.25']
/cvs/cds/caltech/apps/linux64/python/lib64/python2.4/site-packages/nds/__init__.py:28: RuntimeWarning: No protocol specified, attempting protocol nds_v2
  super(daq, self).__init__(host, port)
Connecting NDS2 .... authenticate done
Traceback (most recent call last):
  File "./setSensors.py", line 81, in ?
    mean = acquire(x)
  File "./setSensors.py", line 73, in acquire
    daq.request_channel(chans[x])
Boost.Python.ArgumentError: Python argument types in
    daq.request_channel(daq, str)
did not match C++ signature:
    request_channel(_daq_t {lvalue}, daq_channel_t*)

I'm not exactly sure what the problem is.  Line 73, looks like it should have 2 arguments in the daq.request_channel, but even if I put in the "daq" variable (which is set a few lines above), I get the exact same error.  So...something else is wrong.  Ideas from someone who "speaks" python??

 My guess is that this has something to do with the NDS client version you're using.  Try running the script on a machine where pynds and nds-client are known to be compatible, like pianosa.

 Doesn't work on pianosa either. Has someone changed the python environment?

pianosa:SUS_SUMMARY 0> ./setSensors.py 1000123215 600 0.1 0.25
Traceback (most recent call last):
  File "./setSensors.py", line 2, in <module>
    import nds
ImportError: No module named nds

  5673   Sun Oct 16 02:30:00 2011 ranaUpdateElectronicsTesting REFL165

Unless the bias feedback circuit has been tuned for the 1 mm diode, its possible that you are seeing some C(V) effects. Its easy to check by looking at the phase response at 165 MHz v. the DC photocurrent. Then the feedback or feedforward gain can be tuned.

 

  5672   Sat Oct 15 17:06:20 2011 KojiUpdateLSCInstallation REFL165

REFL165 was installed on the AP table last night.

Meanwhile I found the DC power level at the REFL PDs were 0.8~1.2V if the PRM is aligned and the IFO is not locked.
This corresponds to 16~24mA (20~30mW). This is too big.

The HWP of the REFL path were adjusted so that we have 6~10mA (8~12mW) on each PDs.

  5671   Sat Oct 15 16:42:08 2011 KojiUpdateLSCTesting REFL165

Test results of new REFL165 (the first attachment)

- The resonant freq 166.2MHz, Q=57 (previous Q was ~7)

- If we believe the TF measurement, the transimpedance at the resonance is 7.8k [V/A] and the shotnoise intercept current of ~1mA.
The linearity of the peak was confirmed by changing the modulation level of the beam.

- There is a riddle: the white light test indicates 4.5k [V/A] and 2.8mA for those numbers.
There are big descrepancies from those by the TF measurements.


Further analysis of the descrepancies:

Using the noise measurements with different DC current levels, the transimpedance for each frequency can be reconstructed.

Does this indicate the satiration by the white light???

- The TF measurement shows consistent mag&phase relationship at the resonance (c.f. LISO fit).
So this steep resonance is not an artifact by a noise or glitch but the real structure of the electronics.

- The TF measurement has been done with the photocurrent of ~0.3mA, while the transimpedance measurement
with the white light illumination has the practical effect only when the DC photocurrent is larger than 1mA
because of the circuit noise. Does this higher photo current affected the resonance?

- The off-resonant transimpedance agree with the TF measurement as far as we can see with those measurements.
This may mean that the actual resonant structure has been affected in the white light measurement.
(i.e. not the saturation of the RF opamp which causes the change of the gain at any freq.)
Is the above mentioned higher DC current causing the change of the diode capacitance or other property of the diode or the inductors???

  5670   Sat Oct 15 16:01:26 2011 kiwamuUpdateIOOabout LOCKIN module

Quote from #5669

To make things faster, I think we can just make a LOCKIN which has 3 inputs: it would have one oscillator, but 6 mixers. Should be simple to make.

 I think the idea of having multiple inputs in a LOCKIN module is also good for the LSC sensing matrix measurement.

Because right now I am measuring the responses of multiple sensors one by one while exciting a particular DOF by one oscillator.

Moreover in the LSC case the number of sensors, which we have to measure, is enormous (e.g. REFL11I/Q, REFL33I/Q, REFL55I/Q, ... POY11I/Q,...) and indeed it has been a long-time measurement.

  5669   Sat Oct 15 10:58:32 2011 ranaUpdateIOOMC WFS Output Matrix determination

In order to save time and sanity, you should not measure the pitch ->yaw and yaw-> pitch. It makes things too complicated and so far is just not significant. In the past we do not use these for the matrix work.

i.e. there should just be a 3x3 pitch matrix and a 3x3 yaw matrix. Once the loops are working we could investigate these things, but its really a very fine tweak at the end. There are quite a few other, more significant effects to handle before then.

To make things faster, I think we can just make a LOCKIN which has 3 inputs: it would have one oscillator, but 6 mixers. Should be simple to make.

  5668   Sat Oct 15 04:53:41 2011 SureshUpdateIOOMC WFS Output Matrix determination

After we had a rough idea of what the output matrix looks like (see this elog),
I tried to measure the transfer function coefs (TFCs) between mirror degrees of freedom and the WFS sensors (WFS1, WFS2 and MC_Trans QPD)
I found that the TFCs that I obtained at 10.15 Hz did not have any resemblance to the previously identified output matrix.

The problem, I realised, arises because the various lockins used
in the C1IOO model do not have the same relative phase; So if we try to excite a mirror with one of them
and demodulate a sensor signal on any of the other lockins the resulting output would not have the correct phase
(relative to the 1st lockin output). As a result unless we can reset the phase of all the lockins
simultaneously, we cannot demodulate multiple signals at the same time. (Joe/Jamie, Is it possible to
reset/reinitialise the phase of the CLK signals of the lockings? )


To get around this problem Koji suggested that I use just one lockin and determine all the 36 elements of the transfer matrix with it one at a
time rather than six at a time. When I did that, I got results consistent with the previoulsly determined outmatrix. It, of course, takes six times longer.

The matrix I first got is this one

 

(Mag, Phase) WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 0.332 0.518 0.316 0.019 0.066 0.000
  5.832 1.892 8.180 38.285 8.807 0.000
             
MC2P 0.355 1.798 0.342 0.023 0.144 0.000
  72.977 76.683 76.804 -16.364 77.451 71.579
             
MC3P 0.352 0.394 0.254 0.036 0.023 0.000
  2.005 3.249 6.249 5.712 26.349 NAN
             
MC1Y 0.051 0.055 0.058 0.788 1.024 0.001
  15.979 -4.487 -9.707 2.642 1.276 0.000
             
MC2Y 0.142 0.044 0.130 1.966 0.579 0.017
  70.044 83.818 76.397 74.283 76.134 77.269
             
MC3Y 0.044 0.052 0.022 0.080 0.948 0.194
  22.932 14.227 -45.924 9.677 1.125 1.124
             
Which can be  recast as below          
             
             
Magnitude WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 0.332 0.518 0.316 0.02 0.07 0
MC2P 0.355 1.798 0.342 0.02 0.14 0
MC3P 0.352 0.394 0.254 0.04 0.02 0
MC1Y 0.05 0.05 0.06 0.788 1.024 0.001
MC2Y 0.14 0.04 0.13 1.966 0.579 0.017
MC3Y 0.04 0.05 0.02 0.080 0.948 0.194

Phase WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 5.8 1.9 8.2 38.3 8.8 0.0
MC2P 73.0 76.7 76.8 -16.4 77.5 71.6
MC3P 2.0 3.2 6.2 5.7 26.3 NA
MC1Y 16.0 -4.5 -9.7 2.6 1.3 0.0
MC2Y 70.0 83.8 76.4 74.3 76.1 77.3
MC3Y 22.9 14.2 -45.9 9.7 1.1 1.1

 

Note that when MC2 is excited all the sensors showed a response about 75 deg out of phase with the reference (MC1 --> WFS1_PIT ) This was traced to the fact that while there is a 28Hz Elliptic LP filter on

both MC1 and MC3, while it is absent on MC2.  The Transfer functions  below show the difference in the phase of their response

WFS1P_RespTo_MC1andMC2.pdf

 

Since the MC2 POS is used in servos involving MCL we cannot afford to install a 28 Hz LP filter into the MC2 coil drivers.  However a module with the 28 Hz ELP was switched on, in each of the

 MC2 PIT and YAW filter banks.   I then checked to see if this has affected the relative phase of variour sensors.  The Phase angle between I and Q on each sensor channel was checked and corrected. 

Below are the spectra with the "before" and "after" correction of phases.

Before:

 WFS1_IQphase20111015_1.pdf        WFS2_IQphase20111015_1.pdf

 

Obviously this needed adjustment to reduce Q phase.   

  After twealkng the angle "R":

WFS1_IQphase20111015_2.pdf      WFS2_IQphase20111015_2.pdf

 

And again determined the transfer matrix (below). 

( I , Q ) WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 0.236 -0.300 0.229 0.049 -0.008 0.000
  0.015 -0.004 -0.027 0.011 -0.019 0.000
             
MC2P -0.125 -0.962 -0.135 0.114 0.028 0.000
  0.007 -0.052 -0.028 -0.004 -0.002 0.000
             
MC3P -0.225 -0.254 -0.255 -0.026 -0.010 0.000
  0.004 -0.012 -0.010 0.009 0.002 0.000
             
MC1Y -0.059 -0.023 -0.040 0.460 0.705 0.001
  0.004 0.003 0.009 0.009 0.017 0.000
             
MC2Y 0.030 0.190 0.040 -1.144 -0.296 0.015
  0.007 0.006 -0.009 -0.038 -0.009 0.001
             
MC3Y 0.018 -0.108 -0.018 0.134 -0.832 -0.001
  0.017 0.005 0.001 0.006 -0.016 0.000
 

Magnitude WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 0.236 0.300 0.231 0.05 0.02 0
MC2P 0.125 0.964 0.138 0.11 0.03 0
MC3P 0.225 0.254 0.255 0.03 0.01 0
MC1Y 0.06 0.02 0.04 0.460 0.705 0.001
MC2Y 0.03 0.01 0.19 1.145 0.296 0.015
MC3Y 0.02 0.11 0.02 0.134 0.832 0.001

Phase WFS1P WFS2P MC_T_P WFS1Y WFS2Y MC_T_Y
MC1P 3.694 0.784 -6.778 13.1 66.67 #DIV/0!
MC2P -3.214 3.100 11.557 -2.05 -4.48 0
MC3P -1.020 2.665 2.158 -19.1 -10.76 NA
MC1Y -3.96 -6.45 -12.14 1.085 1.357 0.000
MC2Y 13.22 41.08 -2.6 1.887 1.706 4.987
MC3Y 42.69 -2.56 -3.73 2.652 1.068 0.000

 

This time the signals are all nearly in the same phase and in agreement with the  outmatrix estimate made earlier.

 

I plugged these TFCs into the matrix inversion code: wfsmatrix2.m.   And get the following inverse:

 

  WFS1P_Act WFS2P_Act MC_Trans_P_Act WFS1Y_Act WFS2Y_Act MC_TRANS_Y_Act
MC1P 1 -0.64        
MC2P -0.27 -1        
MC3P 0.98 -0.65        
MC1Y       -0.26 -1  
MC2Y       1 0.12  
MC3Y       0.16 0.07  

 

I have ignored the MC2_Trans_P and Y sensors for now.

  5667   Fri Oct 14 18:38:41 2011 kiwamuUpdateSUSC1:SUS-ETMX_SPDMon fixed

Quote from #5666

 Anyway, things are working now:

 Good job ! Thank you so much

  5666   Fri Oct 14 16:20:11 2011 ZachUpdateSUSC1:SUS-ETMX_SPDMon fixed

I offered to help Kiwamu with some of the 40m work. The first task was to figure out why the ETMX side OSEM monitor was so low, since we know that the depth is about right. It was showing ~0.13 V to the others' ~0.7 V.

TL,WR: There was a wire disconnected from the breakout panel on the side of the rack

I started by pulling the board out and checking to make sure that it was working properly. I injected a sine wave to the SIDE IN and found that it showed up in the signal coming out of the back (into the crate) just fine (see below). One strange thing I noticed while testing the board is that both inputs for each used channel of the MAX333 switches on the board are shorted to their respective outputs. That is, the switches seem to be open to BOTH 0 and 1 logic states. This seems counterintuitive, but perhaps there's something about how these work that I don't know.

board_works.png

 

Then I went about tracing the signal from the back of the crate to the breakout panel on the side of the rack. I opened it up, verified that the ribbon cable was transmitting correctly, and as I went to plug it back in I noticed that one of the wires---the correct one---had come completely undone.

rut_roh_raggie.png

The screw clamp appeared to be a bit finicky, as I had to loosen and tighten it a few times before it finally seemed to grab hold of the wire. It probably just wasn't tight in the first place and the wire was pulled out. Anyway, things are working now:

Screen_shot_2011-10-14_at_4.09.45_PM.png

  5665   Fri Oct 14 04:35:45 2011 kiwamuUpdateLSClocking tonight

The lock of DRMI wasn't stable enough to measure the sensing matrix. Failed.

PRMI and SRMI were okay and in fact they could stay locked robustly for a long time.

I added a new option in the C1IFO_CONFIGURE screen so that one can choose Signal-Recycled Michelson in carrier resonant condition.

Screen_shot_2011-10-14_at_4.24.01.png

 Additionally the orthogonalization of the I-Q signals on REFL55 should be done because it hasn't been done.

 

  5664   Thu Oct 13 23:58:38 2011 KojiUpdateLSCfixing REFL165

I already have reported in this entry that REFL165 shows too high DC output which does not depend on the light level on the diode.
Today I removed REFL165 from the table and inspected it.

The diode has been burnt as shown in the first picture (left).
The window is smoked, and the photo sensitive surface has been removed from its base. It moves in the can.

The burnt diode was replaced to the new one.
The new one shows ~30% better capacitance of ~50pF
and I had to increase the inductance from 14nH (i.e. 15nH//220nH) to 18nH.
After some struggles to increase/decrease the stray inductance by moving the SMD capacitors a little, the resonance is reasonably tuned to 166MHz.

The comprehensive test will be performed shortly.

  5663   Thu Oct 13 21:44:48 2011 MirkoUpdateCDSSeismic BLRMS channels, new RMS calculation

[Rana, Koji, Mirko]

We looked into the CDS RMS block c-code as described in Rolfs RCG app guide. Seems the block uses a first order LP filter with a corner freq. / time of 20k execution cycles. There are also some weird thersholds at +-2000counts in there.

I was looking into implementing a hand-made RMS block, by squaring, filtering, rooting. The new RMS (left) seems nicer than the old one (bottom right). Signal was 141counts sinus at 4Hz.

Filters used: Before squaring: 4th order butterworth BP at given freq. & (new) 6th order inverse Chebyshew 20dB at 0.9*lower BP freq. and 1.1*upper BP freq. => about 1dB at BP freq.

                       After squaring: 4th order butterworth LP @ 1Hz.

C1PEM execution time increased from about 20us to about 45us.

Made a new medm screen with the respective filters in place of the empty C1PEM_OVERVIEW. Should go onto the sitemap.

New_RMS_vs_old_RMS.png

Original RMS LP is slower than 0.1Hz, see below for single LP at 0.1Hz in the new RMS. Original RMS is faster than single LP @ 0.01Hz

Original_RMS_LP_slower_than_0.1Hz.png

Some of the channels are recorded as 256Hz DAQ channels now. Need to figure out how to record these as 16Hz EPICS channls.

  5662   Thu Oct 13 21:40:59 2011 ranaSummaryVACRecovery from the power shutdown is completed

 As it turns out Steve is not crazy in this particular instance: the vacuum computer, linux3 , has some issues. I can login just fine, but trying to open a terminal makes the CPU rail and the RAM max out and eventually the machine freezes.

Under KDE, I can open a terminal OK as root, but if I then try a 'su controls', the same issue happens. Let's wait for Jamie to fix this.

  5661   Thu Oct 13 20:25:32 2011 KatrinUpdateGreen LockingLPF transfer function YARM

It is a 4th order filter with cut of frequency of 120 kHz.

 

Design

designLPF.png

 

Measurement

 LC_LPF.TIF

  5660   Thu Oct 13 14:23:09 2011 steveUpdateSUSITMX oplev with 3 mm beam on qpd

 I replaced the JDSU-Uniphase 1125P by 1103P He/Ne laser. This new laser had 2.8 mW output yesterday. It degraded to 0.5 mW by this morning.

The beam size on the QPD is ~3 mm  This should give us  better sensitivity. These are not the perfect lenses at all, but we have them here.

On the other hand, there are still some coherence below 1 Hz, so the laser intensity noise or clipping dominating  this  part of the spectrum.

 

  5659   Thu Oct 13 03:22:53 2011 kiwamuUpdateLSCmeasurement of sensing matrix : just began

- status update on LSC activity :

The measurement of the LSC sensing matrix has begun. But no useful results yet.

 

 The measurement script (#4850) ran pretty well after I did some modifications to adopt the script to the latest LSC model.

However the SNR weren't so great particularly in REFL33 in the PRMI configuration.

So I will tune the amplitude of excitations and integration times tomorrow.

Currently the excitation is at 238.1 Hz, where no disturbing structures are found in the spectra.

  5658   Wed Oct 12 19:58:32 2011 KatrinUpdateGreen LockingPower splitter is unbalanced

The mini circuit power splitter ZFRSC-42S+ used at the YARM has no balanced output as it should have according to the data sheet.

@ 0.05MHz  the amplitude unbalance should be 0.03 dB

A quick measurement shows that there is a LO amplitude dependent unbalance:

LO amplitude input (Vpp)  unbalanced output (dB)
1.3 3.66
1.4 4.08
1.5 4.28
1.6 4.36

So my question is, shall I replace the power splitter just in case it is further degrading?

  5657   Wed Oct 12 18:54:02 2011 KatrinUpdateGreen Locking60 Hz oscillation due to broken BNC cable

There was a 60 Hz and 120 Hz oscillation on the green PDH photo diode output. After a long search, I could identify that

the source was a broken BNC cable which was connected to the photo diode. I exchanged that BNC cable and the 60 Hz

and 120 Hz are gone :-)

With the new cable the PD output was less noisy so that it was easier to achieve a better alignment of the light to the cavity.

The reflected power could be reduced from 40% to 30%. For perfect alignment the reflected power would be 20%.

  5656   Wed Oct 12 17:53:01 2011 kiwamuUpdateLSCBS actuator response : fitting done and histroy of delays

An update on calibration of the BS actuator : A fitting has been done.

(Fitting)

 I used LISO for fitting the complex transfer function.
Because the data points around 1 Hz didn't have big coherence a few data points, which had coherence of less than 0.9, were excluded.
Also the fitting of the Q-factor wasn't successful due to the lack of good data points around the resonance.So I left Q fixed to be 5 in the fitting.
 

(Fitting result)

G =  2.18060874008e-8 +/- 6.425e-10 (2.95%)
f0 =  1.0100491195 +/- 1.51e-2 (1.49%)  [Hz]
Q = 5 (fixed)
delay =  423.2753462089e-6 +/- 4.989e-6 (1.18%)  [sec]

bode_BSactuator.png

 

(History of delay)
 Because we have been observing several different amount of delays in different configurations, perhaps it is worth to summarize those numbers.
     description  delay [usec]   elog entry
       MICH lock (BS actuation)  423 this entry
       LSC feed forward path  127 #5218
       MICH lock (BS actuation)  600 #4638
      ALS on X arm (ETMX actuation)  330 #4196
      RFM (from c1lsc to c1sus)  125 #4153
      from ADC to DAC (all the front end machine)  38-110 #3961
      from ADC to DAC (c1sus)  124 #3838
     RFM (c1ioo and c1sus)   8-62 #3855

Quote from #5648
Tomorrow I will do a serious fitting.

ELOG V3.1.3-